OpenStudy (anonymous):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. A.4 1/20 B.1 7/20 C.2/3 D.9/20

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@izzy2917

OpenStudy (anonymous):

@omgitsjc

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

@BSwan

OpenStudy (anonymous):

3/5 + 3/4 = 12/20 + 15/20 = 27/20 (total juice) 3/(27/20) = 1/ (x) --- 3 servings to 27/20 juice = 1 serving to x juice cross multiply (3)(x) = (27/20)(1) 3x = 27/20 -- divide by 3 x = (27/20) / 3 x = 27/20 * 1/3 x = 27/60 which reduces to 9/20

OpenStudy (anonymous):

so your answer is 9/20 ;)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no prob!